OpenStudy (anonymous):

find weather y varies with x. if so find the constant of variation k. –6y = –5x

OpenStudy (anonymous):

@loopki40

OpenStudy (anonymous):

@iGreen

OpenStudy (dan815):

rearrange in terms of y=...

OpenStudy (dan815):

divide both side by -6 -6/-6 *y=-5/-6 * x y=5/6 x

OpenStudy (dan815):

therefore K = 5/6